### v1

# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

Abstract
A framework for portfolio allocation based on multiple hypotheses prediction using structured ensemble models is presented. Portfolio optimization is formulated as an ensemble learning problem, where each predictor focuses on a specific asset or hypothesis. The portfolio weights are determined by optimizing the ensemble’s parameters, using an equal-weighted portfolio as the target, serving as a canonical basis for the hypotheses. Diversity in learning among predictors is parametrically controlled, and their predictions form a structured input for the ensemble optimization model. The proposed methodology establishes a link between this source of learning diversity and portfolio risk diversification, enabling parametric control of portfolio diversification prior to the decision-making process. Moreover, the methodology demonstrates that the diversity in asset or hypothesis selection, based on predictions of future returns, before and independently of the ensemble learning stage, also contributes to the out-of-sample portfolio diversification. The sets of assets with more diverse but lower average return predictions are preferred over less diverse selections. The methodology enables parametric control of diversity in both the asset selection and learning stages, providing users with significant control over out-of-sample portfolio diversification prior to decision-making. Experiments validate the hypotheses across one-step and multi-step decisions for all parameter configurations and the structured model variants using equity portfolios.
keywords:
diversity , ensemble learning , multiple hypotheses prediction , portfolio optimization , predict-then-optimize , structured models
journal: Expert Systems with Applications
[…]
[inst1] organization=University of Reading, Department of Computer Science, addressline=Whiteknights House, city=Reading, postcode=RG6 6UR, country=United Kingdom, email=a.j.rodriguezdominguez@pgr.reading.ac.uk
1 Introduction
The problem of portfolio allocation with a quantitative approach was introduced by Harry Markowitz in 1952 in the so-called Modern Portfolio Theory (MPT) (Markowitz, 1952). It consists of two phases: asset selection and portfolio weight optimization. The asset selection phase decides which assets from a larger set, such as a market, will be included in the portfolio, which has a fixed number of assets but not predetermined choices or weights. The portfolio weight optimization phase determines the percentage of capital allocated to each selected asset. It also has the merit of introducing the concept of diversification as a fundamental component in risk management for investments. In this specific case, diversification is induced by the covariance matrix (Markowitz, 1952). Since then, a large number of methods known as ”plug-in” (or data-driven) approaches have been developed which use optimization inputs with improved estimates of future returns (Best and Grauer, 1991; Broadie, 1993; Giglio et al., 2022).
The standard data-driven approach to portfolio choice optimization replaces population parameters with their sample estimates, which leads to poor out-of-sample performance due to parameter uncertainty (Zhang et al., 2018). This issue is typically addressed by deriving expected loss functions that quantify the impact of using sample means and covariance matrices to estimate the optimal portfolio (Kan and Zhou, 2007). Other researchers have focused on portfolio optimization using robust estimators, which can be computed by solving a single non-linear programming problem (DeMiguel and Nogales, 2007; Lassance et al., 2023). Shrinkage techniques adjust extreme coefficients toward more central values in the sample covariance matrix to reduce estimation error, often incorporating a parameter to control the level of shrinkage (Ledoit and Wolf, 2003; Bodnar et al., 2024).
Few researchers have also employed the use of predictive models to optimize decision-making processes, often with the objective of minimizing decision-related costs. Such conventional ”predict-then-optimize” paradigm decouples prediction and decision-making into two distinct stages: first, a predictive model is developed to maximize predictive accuracy; second, decisions are derived based on the model’s outputs and associated cost functions. A significant limitation of this approach is its failure to incorporate cost considerations during the predictive modeling phase (Donti and Kolter, 2017). To address this shortcoming, recent advances have introduced integrated methodologies such as ”predict-and-optimize” (Wilder et al., 2019; Vanderschueren et al., 2022) and Smart ”Predict, then Optimize” (SPO) (Elmachtoub and Grigas, 2017). These approaches bridges the gap between structured prediction and cost-sensitive optimization, offering a robust solution for enhancing decision-making processes (Kotary et al., 2023). These approaches, however, are computationally expensive and mostly focus on solving the classification problems (Goh and Jaillet, 2016).
While the aforementioned solutions reduce parameter uncertainty and can improve out-of-sample performance, they fail to establish a connection between in-sample and out-of-sample diversification, which is a key feature addressed in this work. In this context, a framework is proposed for portfolio optimization using structured ensembles for prediction with multiple hypotheses. Each portfolio component corresponds to a hypothesis, with its return series modeled by an individual predictor. This approach differs from return forecasting (Ma et al., 2021) or scoring models (Nguyen and Lo, 2012; Zhao et al., 2022), where models are used to score assets and predict their returns before applying optimization techniques for portfolio allocation. Building on prior research, the degree of diversity in ensemble learning is parametrically controlled during predictor training and encoded in a structured data set of predictions (Dominguez et al., 2025). This data set is used as input for an ensemble model, with the equal-weight portfolio serving as the target, thereby creating a canonical hypothesis space for portfolio predictions. Scaled ensemble parameters subsequently determine portfolio weights in the allocation problem. To the best of the authors’ knowledge, this is the first plug-in method enabling parametrically controlled portfolio risk diversification prior to the optimization step.
Additionally, this work also investigates the hypothesis that out-of-sample portfolio diversification relates to the asset or hypothesis selection. Diversity in asset selection is controlled parametrically, independently of the portfolio optimization phase. A parameter regulates the diversity of asset sets—randomly selected from ranked out-of-sample return predictions—across various portfolio sizes and parameters. Results show that greater diversity in asset sets increases out-of-sample portfolio diversification and risk-adjusted returns, even when the sets have lower average returns. This work links both sources of diversity (i.e., asset selection and portfolio weight optimization) to out-of-sample portfolio diversification and demonstrates that the diversification levels from both sources can be parametrically controlled prior to decision-making, offering valuable insights for practitioners.
Following are the main contributions proposed in this work:
  • 1.
    Introduced a consistent framework for using structured ensemble models in multiple hypothesis prediction settings for portfolio optimization, facilitated by the equal-weighted portfolio target selection.
  • 2.
    Presented a plug-in method with parametric a priori control over the portfolio’s out-of-sample risk diversification level.
  • 3.
    Proposed a strategy to parametrically control the diversity linked to asset or hypothesis selection with out-of-sample portfolio diversification.
  • 4.
    Systematically validated the link between diversity in structured ensembles for multiple hypothesis prediction and risk diversification across all experiments and models, for both one-step and multi-step ahead decision cases.
The paper is organized as follows. Section 2 provides a review of previous work on data-driven and robust optimization, Predict-and-Optimize frameworks, structured ensemble learning, and diversity. Section 3 describes the proposed framework and methodology. Section 4 presents the experimental results and discussions. Finally, Section 5 offers the concluding remarks and future outlook.
2 Literature Review
2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
2.2 Predictive Modeling Integrated with Decision Optimization
Predictive models have become essential tools for optimizing decision-making processes, particularly when the objective is to minimize decision-related costs. However, this approach overlooks cost considerations during the predictive modeling phase, leading to suboptimal decision quality (Donti and Kolter, 2017). Recent advancements have sought to address this issue through integrated frameworks like ”predict-and-optimize” (PO), Smart Predict-then-Optimize (SPO) and End-to-End Predict-Then-Optimize (EPO) (Elmachtoub and Grigas, 2017, 2022; Vanderschueren et al., 2022). These frameworks use surrogate loss functions to manage computational challenges associated with original loss functions and link structured prediction outputs to decision variables (Elmachtoub and Grigas, 2017). The SPO+ framework further advances this integration by defining loss functions based on the objective costs of nominal optimization problems, improving computational efficiency and decision-aware modeling (Elmachtoub and Grigas, 2022). However, most existing studies in structured prediction focus on classification problems (Goh and Jaillet, 2016) and are computationally intensive, as they require backpropagation through optimization solutions, which often involves handcrafted rules, and they face challenges in handling nonconvex or discrete optimization models (Kotary et al., 2023).
2.3 Diversity in Ensemble Learning and Portfolio Diversification
A unified theory of ensemble diversity explains that diversity is a hidden dimension in the bias-variance decomposition of the ensemble loss (Wood et al., 2024). Moreover, studies on diversity measures in ensemble learning show their significant impact on model accuracy, generalization, and robustness (Ortega et al., 2022).
Researchers link portfolio estimation to statistical decision theory by treating the difference in investor utility between the ”true” optimal portfolio and the plug-in portfolio as an economic loss function (Kan and Zhou, 2007). Outcomes are enhanced by incorporating the investor’s utility objective directly into the statistical weight estimation process like in Bayesian decision theoretic approaches, rather than addressing estimation and utility maximization as separate issues (Avramov and Zhou, 2010; Kan and Zhou, 2007). Research has demonstrated the economic implications of using Mean Square Error (MSE) in portfolio optimization and financial forecasting, showing that integrating MSE into decision-making processes can enhance portfolio performance by minimizing estimation errors and optimizing risk-return trade-offs (Lasse Heje Pedersen and Levine, 2021; Mahadi et al., 2022; Cai et al., 2024). This work extends this idea by incorporating diversity into the MSE loss function within an ensemble learning framework. Furthermore, it establishes for the first time a connection between this notion of diversity and portfolio diversification, proposing parametric methods to manage diversification prior to decision-making.
3 Framework Description
3.1 Preliminary
Let $\boldsymbol{\mu}=\begin{bmatrix}\mathbb{E}[f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x})],&\mathbb{E}[f_{\boldsymbol{\theta}_{2}}(\boldsymbol{x})],&\dots&\mathbb{E}[f_{\boldsymbol{\theta}_{N}}(\boldsymbol{x})]\end{bmatrix}^{\top}$ be the vector of expected returns of the base learners, $\boldsymbol{\Sigma}$ be the covariance matrix where $\Sigma_{ij}=\text{Cov}(f_{\boldsymbol{\theta}_{i}}(\boldsymbol{x}),f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}))$, and $R_{T}$ be the target return for the portfolio. Let $\boldsymbol{\hat{w}}$ be an estimator of $\boldsymbol{w}$ based on an i.i.d. sample from a distribution $\mathcal{X}$. The ensemble prediction error can be decomposed into bias and variance components. In this case, the bias is minimized subject to the portfolio return being above the target value, and the variance is minimized. The optimization problem can be formulated as follows.
$\min_{\boldsymbol{w}}\quad\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}+\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ (1)
$\quad\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$
$\quad\boldsymbol{w}^{\top}\boldsymbol{1}=1$
where $\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}$ represents the bias of the expected return of the portfolio compared to the target return, and $\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ represents the variance of the portfolio. $\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$ ensures that the expected return of the portfolio is above the target return. This setup allows to find the optimal weights $\boldsymbol{\hat{w}}$ that minimize the total prediction error of the portfolio, considering both bias and variance, while ensuring that the expected return of the portfolio is above the target value. The expression in (1) generalizes the standard MSE (when $A=I$), reflecting the bias-variance trade-off (Cai et al., 2024):
$MSE_{A}(\boldsymbol{\hat{w}})=\mathcal{L}_{A}(\mathbb{E}[\boldsymbol{\hat{w}}]-\boldsymbol{w})+\text{tr}(A\text{Var}[\boldsymbol{\hat{w}}])$ (2)
where the first term vanishes for an unbiased estimator, and the second term vanishes if the estimator has zero variance. Now, consider $\boldsymbol{w}$ as the solution to a decision problem:
$\boldsymbol{w}=\arg\max_{\boldsymbol{w}\in C}P(\boldsymbol{w})$ (3)
for some performance measure $P$ and constraint set $C$. The decision loss for $\boldsymbol{\hat{w}}\in C$ is
$D_{P}(\boldsymbol{\hat{w}})=P(\boldsymbol{w})-\mathbb{E}[P(\boldsymbol{\hat{w}})]\geq 0,$ (4)
where the expectation is taken under $D$. For a given $\boldsymbol{\hat{w}}$ and matrix $A$, establish a relationship between $MSE_{A}(\boldsymbol{\hat{w}})$ and $D_{P}(\boldsymbol{\hat{w}})$, expressed as a function $f$:
$MSE_{A}(\boldsymbol{\hat{w}})=f(D_{P}(\boldsymbol{\hat{w}}))$ (5)
For a parameterized estimator $\boldsymbol{\hat{w}}(\delta)$, compare $\delta_{MSE}=\arg\min_{\delta}MSE_{A}(\boldsymbol{\hat{w}}(\delta))$ with $\delta_{D}=\arg\min_{\delta}D_{P}(\boldsymbol{\hat{w}}(\delta))$ and assess the decision loss increase:
$\Delta_{P}(\boldsymbol{\hat{w}}(\delta))=D_{P}(\boldsymbol{\hat{w}}(\delta_{MSE}))-D_{P}(\boldsymbol{\hat{w}}(\delta_{D}))=\mathbb{E}[P(\boldsymbol{\hat{w}}(\delta_{D}))]-\mathbb{E}[P(\boldsymbol{\hat{w}}(\delta_{MSE}))]\geq 0$ (6)
If $f$ increases and $\Delta_{P}$ is small, the MSE serves as an effective loss function for the decision problem, although these conditions do not imply each other.
In this work, structured ensemble multi-hypothesis prediction models are applied to portfolio optimization. Next, the Portfolio Structured Ensemble Model (PSEM) is introduced, with a loss function based on the generalized bias-variance-diversity decomposition for ensembles (Wood et al., 2024).
3.1.1 Diversity in Ensemble Learning
In this section, the Bias-Variance-Diversity decomposition (Wood et al., 2024), is formulated and applied to the problem of portfolio optimization using multiple hypothesis structured prediction models.
If a training set $\{(\boldsymbol{R}_{i},R_{Ti})\}_{i=1}^{N}$ drawn from a random variable $D\sim P(\boldsymbol{R},\boldsymbol{R_{T}})^{N}$, with $\boldsymbol{R}_{i}=\{R_{\bullet j}\}_{j=1}^{M}$, and $\boldsymbol{R}=\left[\begin{matrix}R_{11}&\ldots&R_{1M}\\ \vdots&\ddots&\vdots\\ R_{N1}&\ldots&R_{NM}\\ \end{matrix}\right]$. From this, a portfolio ensemble model is learned with predictions, $\boldsymbol{\hat{R}_{T}}$. The expected squared risk can be decomposed as follows:
$\mathbb{E}_{D}\left[\mathbb{E}_{\hat{R}_{T},R_{T}}\left[(\boldsymbol{R_{T}}-\boldsymbol{\hat{R}_{T}})^{2}\right]\right]=$
$\underbrace{\mathbb{E}_{R}\left[\mathbb{E}_{R_{T}|\hat{R}_{T}}\left[(\boldsymbol{R_{T}}-\boldsymbol{R_{T}}^{*})^{2}\right]\right]}_{\text{Noise}}+\underbrace{\mathbb{E}_{\hat{R}_{T}}\left[(\mathbb{E}_{D}[\boldsymbol{\hat{R}_{T}}]-\boldsymbol{R_{T}}^{*})^{2}\right]}_{\text{Bias}}+$
$\underbrace{\mathbb{E}_{\hat{R}_{T}}\left[\mathbb{E}_{D}\left[(\boldsymbol{\hat{R}_{T}}-\mathbb{E}_{D}[\boldsymbol{\hat{R}_{T}}])^{2}\right]\right]}_{\text{Variance}}$ (7)
In this context, $\boldsymbol{R_{T}}^{*}=\mathbb{E}_{R_{T}|R}[\boldsymbol{R_{T}}]$ represent the Bayes-optimal predictions. The first term accounts for the irreducible noise in the problem, which is independent of any model parameters. The second term is the bias, defined as the loss of the expected model relative to the Bayes-optimal predictions. Finally, the third term is the variance, which captures the variation in $\boldsymbol{\hat{R}_{T}}$ due to different training sets.
If $\boldsymbol{R}_{j}$ are the predictions of the $j$-th model in the portfolio ensemble, $\overset{\circ}{\boldsymbol{R}}_{j}$ is the centroid prediction of the $j$-th model in the portfolio ensemble. Assuming a bias-variance decomposition holds in the form of (7), the Generalized Bias-Variance-Diversity decomposition also holds (Wood et al., 2024):
$\mathbb{E}_{D}\left[\mathbb{E}_{\hat{R}_{T},R_{T}}[\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{\hat{R}_{T}})]\right]=\mathbb{E}_{\hat{R}_{T}}\left[\mathbb{E}_{R_{T}|\hat{R}_{T}}[\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R_{T}}^{*})]\right]+\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}}^{*},\overset{\circ}{\boldsymbol{R}}_{j})+$ (8)
$\quad+\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{D}\left[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})\right]-\mathbb{E}_{D}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}_{T}},\boldsymbol{R}_{j})\right]$
where $\overset{\circ}{\boldsymbol{R}}_{j}=\arg\min\mathbb{E}_{D}[\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})]$ and the combiner is
$\boldsymbol{\hat{R}_{T}}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})$, equivalent to the equal-weighted-portfolio ($1/N$-portfolio rule (DeMiguel et al., 2009)). This decomposition separates the expected ensemble risk into: the noise, which represents the irreducible error inherent in the problem; the average bias of the portfolio returns predictions $\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\overset{\circ}{\boldsymbol{R}}_{j})$; the average variance of the individual models’ predictions in the portfolio ensemble $\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})]$; the expected ambiguity term, considered as the ensemble diversity, which has the opposite sign to the other terms, $\mathbb{E}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}_{T}},\boldsymbol{R}_{j})\right]$, capturing the disagreement among the portfolio ensemble members in their predictions.
The Bias-Variance-Diversity decomposition using the MSE, $\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{\hat{R}_{T}})=(\boldsymbol{R_{T}}-\boldsymbol{\hat{R}_{T}})^{2}$, which implies $\boldsymbol{\hat{R}_{T}}=\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{R}_{j}$, and $\overset{\circ}{\boldsymbol{{R}_{T}}}=\mathbb{E}_{D}[\boldsymbol{\hat{R}_{T}}]$, at each data set $(\boldsymbol{R},\boldsymbol{R_{T}})$ is:
$\mathbb{E}_{D}\left[(\boldsymbol{\hat{R}_{T}}-\boldsymbol{R_{T}})^{2}\right]=$
$=\underbrace{\frac{1}{M}\sum_{j=1}^{M}(\overset{\circ}{\boldsymbol{{R}_{j}}}-\boldsymbol{R_{T}})^{2}}_{\text{Average Bias}}+\underbrace{\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{D}\left[(\boldsymbol{{R}_{j}}-\overset{\circ}{\boldsymbol{{R}_{j}}})^{2}\right]}_{\text{Average Variance}}-\underbrace{\mathbb{E}_{D}\left[\frac{1}{M}\sum_{j=1}^{M}(\boldsymbol{{R}_{j}}-\boldsymbol{\hat{R}_{T}})^{2}\right]}_{\text{Diversity}}$ (9)
3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
In this portfolio structured ensemble model framework, the goal is to make portfolio allocation decisions at time $t$ for an investment horizon $\Delta t=T-t$. The daily time-series returns of a portfolio for the period $\Delta t$ are denoted as $\boldsymbol{R_{T}}=\{R_{Ti}\}_{i=t}^{T}$. To approximate the expected return, a structured ensemble model predicts the portfolio time-series returns as $\boldsymbol{\hat{R}_{T}}$. This model consists of $M$ individual predictors, $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, where each predictor focuses on forecasting the return of the $j$-th component of the portfolio $\{\boldsymbol{R}_{j}\}_{j=1}^{M}$. The predictors are trained using stochastic gradient descent (SGD). During each training iteration $i$, for a given input $\{\boldsymbol{x}_{ij}\}_{i=1,j=1}^{N,M}$ and its corresponding target $R_{ij}$, the parameters of the $j$-th predictor are updated as:
$\boldsymbol{\theta}_{j}=\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}),R_{ij})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta\left(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))\right),$
where $\eta_{j}$ is the learning rate, and the loss function $\mathcal{L}$ combines a squared error term with L2-regularization and parameter $\lambda_{p}$,
$\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}),R_{ij})=\|f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})-R_{ij}\|_{2}^{2}+\frac{\lambda_{p}}{2N}\sum_{j=1}^{M}\|\boldsymbol{\theta}_{j}\|_{2}^{2}$. The indicator function $\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})))$, which adjusts the extent to which each predictor’s parameters are updated, is defined as:
$\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})))=\begin{cases}1-\varepsilon,&\text{if }R_{ij}\in\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$
Here, $\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))$ indicates whether the $j$-th predictor’s output is the top prediction for the $i$-th instance. A smaller diversity parameter $\varepsilon\in[0,1]$, places more emphasis on the top-performing predictors, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble. Once training is complete, the predictors can be used to generate a structured data set from the initial set of inputs, or the predictions obtained during training can be used directly as the structured data set. The predictions of all predictors are organized into a structured data set:
$\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)=\begin{bmatrix}f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{11})&\cdots&f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{1M})\\ \vdots&\ddots&\vdots\\ f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{N1})&\cdots&f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{NM})\end{bmatrix}$ (10)
This structured data set serves as input to a structured ensemble model for prediction. The ensemble output is computed as:
$\boldsymbol{\hat{R}_{T}}=\mathcal{F}(\boldsymbol{\hat{R}}(\varepsilon))\boldsymbol{w}$ (11)
where $\mathcal{F}$ is a set of specific functions of the ensemble model, $\boldsymbol{w}=\{w_{j}\}_{j=1}^{M}$ represents the parameters of the ensemble, which, when appropriately scaled, are equivalent to the portfolio weights, and $\boldsymbol{\hat{R}_{T}}$ represents the predicted time series returns for the period $\Delta t$. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
Figure 1 shows a sketch of the model architecture, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $(\varepsilon)$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the s-RBFN shown in this figure and trained using least squares, described in detail in Section 3.3 (Dominguez et al., 2025).
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3 (Dominguez et al., 2025). $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{i})$ is the $j^{th}$ shooter (predictor), $\hat{R}_{ij}$ is the $i^{th}$ shoot of $j^{th}$ shooter and $\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{R_{T}}$ is the portfolio target, and $\boldsymbol{w}$ is the s-RBFN model parameters optimized by least-squares.
3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
The structured data set is used as input for a radial basis function network, with each $j^{th}$ predictor or hypothesis $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})=\boldsymbol{R}_{j}(\varepsilon)$ associated with a particular basis function $\phi\left(\boldsymbol{R}_{j}(\varepsilon),\mu_{j},\sigma_{j}\right)$, that is, a map $\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right):\mathbb{R}^{N\times M}\to\mathbb{R}^{N\times M}$ is obtained by applying the basis function $\phi\left(\cdot\right)$ to each element of $\boldsymbol{R}(\varepsilon)$. For example, the Gaussian basis function $\phi\left(\cdot\right)=\exp{\left(\frac{-1}{2\sigma_{j}^{2}}\left|{R}_{ij}(\varepsilon)-\mu_{j}\right|^{2}\right)}$ has been used where the parameters of the centers $\mu_{j}$ and the scales $\sigma_{j}$ for the basis functions are calculated from each column of $\boldsymbol{R}(\varepsilon)$ as the mean and standard deviation, respectively. The s-RBFN formulation can now be expressed in matrix form as follows:
$\boldsymbol{\hat{R}_{T}}=\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)\boldsymbol{w}=$
$=\left[\begin{matrix}\phi\left({R}_{11}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left({R}_{1M}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \vdots&\ddots&\vdots\\ \phi\left({R}_{N1}(\varepsilon),\mu_{1},\sigma_{1}\right)&\ldots&\phi\left({R}_{NM}(\varepsilon),\mu_{M},\sigma_{M}\right)\\ \end{matrix}\right]\left[\begin{matrix}w_{1}\\ \vdots\\ w_{M}\\ \end{matrix}\right]$ (12)
The optimal s-RBFN parameters can now be obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{y}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{y}$ (13)
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, including ensuring they sum to one and remain within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\mathbf{w}=1$ and $\mathbf{0}\leq\mathbf{w}\leq\mathbf{1}$.
The time series returns of the target portfolio used to train the structured portfolio ensemble and compute the portfolio weights have the expression $\boldsymbol{R_{T}}=\boldsymbol{w}^{T}\boldsymbol{R}(\varepsilon)$. If (13) is reduced to:
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{R_{T}}=\boldsymbol{P}\boldsymbol{w}^{T}\boldsymbol{x}$ (14)
with $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{R}(\varepsilon)\right)^{\rm T}$, and $\boldsymbol{w}$ assumed to be optimal. In this case, the shape of the constituents’ hypothesis space and the degree of diversification in the structured data set of predictions have already been incorporated into $\boldsymbol{P}$. To ensure that this information is utilized in the optimization of the s-RBFN parameters (portfolio ensemble model weights from (13)), regardless of the terms $\boldsymbol{w}^{T}\boldsymbol{x}$ in (14) for the target portfolio, normalization is applied so that an equally weighted target portfolio $\boldsymbol{R_{T}}$ is assumed, with weights $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the s-RBFN parameters obtained during training are:
$\boldsymbol{\hat{w}}\approx\boldsymbol{w}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (15)
The Centroid Combiner Rule, derived from the Bias-Variance-Diversity decomposition in (8) and adapted to this problem,
$\boldsymbol{\hat{R}_{T}}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R_{T}},\boldsymbol{R}_{j})$, represents the naive portfolio diversification rule (the 1/N portfolio) (DeMiguel et al., 2009).
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from Modern Portfolio Theory (Markowitz, 1952), shown on the left, and a structured Radial Basis Function Network (s-RBFN) optimized via least squares (Dominguez et al., 2025), shown on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
In the mean-variance framework, the key plug-in component is the covariance matrix, which is based exclusively on the input data set $\boldsymbol{x}$ and does not provide a guarantee of achieving portfolio diversification outside the sample. In contrast, the s-RBFN framework incorporates portfolio diversification directly into the modeling process by treating assets as hypotheses and the portfolio as an ensemble predictor, leveraging the diversity among individual predictors during training. The structured data set $\boldsymbol{R}(\varepsilon)$, constructed from predictions generated by models using the same input data as the mean-variance model, enables the use of a continuous diversification parameter $10<\varepsilon<$ 1. This allows for the creation of multiple diversified plug-in data sets from the same input and for portfolio diversification to be adjusted prior to weight optimization.
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the mean-variance framework, least squares optimization can be applied to the input matrix $\boldsymbol{X}$ to determine portfolio weights. In the s-RBFN framework, this input matrix is replaced with the diversified prediction matrix $\boldsymbol{R}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
Figure 2: Systems representation comparison of the Mean-Variance approach to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, denoted as $\boldsymbol{x}=\{\boldsymbol{x_{j}}\}_{j=1}^{M}$, and the target portfolio returns, denoted as $\boldsymbol{y}$, are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{R}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x_{j}})\}_{j=1}^{M}$, parameterized by $\Theta=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, and the s-RBFN model basis functions $\boldsymbol{\Phi}$.
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the s-RBFN method operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{y}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{R}(\varepsilon)$, thereby avoiding any dependence on a specified target.
3.4 Parametric Control of Out-of-Sample Diversification at the Asset or Hypothesis Selection Stage
The portfolio allocation is structured into two stages (Markowitz, 1952): asset selection and portfolio optimization. In the asset selection, the cumulative 1-month forecasts for the 500 members of a hypothetical market are computed and ranked based on their predicted cumulative returns. A threshold $T$ is applied to the ranking, and the top-performing assets above this threshold should be selected based on rational investor behavior. To incorporate diversity into the asset selection process, a parameter $\gamma$ is introduced. This parameter acts as a multiplier of the number of portfolio constituents $M$, generating a sample of size $m=\gamma\times M$ consisting of assets with predictions exceeding the threshold $T$, from which a set of $M$ portfolio constituents is randomly selected.
This can be seen in Figure 3, where the forecast is first computed (left block). Then, depending on whether $\gamma=1$, no diversity is added; if $\gamma>1$, diversity is introduced into the asset selection, provided the sample size is such that all predictions in the sample exceed $T$. From this sample, $M$ portfolio constituents are randomly selected, with the parameter $\gamma$ controlling the degree of diversity in asset or hypothesis selection (central block). The right block refers to the structured prediction model with multiple hypotheses, where the hypotheses have already been selected. At this stage, the parameter $\varepsilon$ regulates the diversity in the learning process of individual predictors and the ensemble. Once the prediction data set with the selected diversity is obtained, the portfolio weights are optimized as described in Section 3.3 (right block in Figure 3). This allows testing the hypothesis that there is a trade-off between the diversity of predicted future returns and their average value as a group. A more diverse group of predictions is preferred, even if it includes lower returns that reduce the group average, highlighting a balance between diversity and the quality of return predictions for asset or hypothesis selection.
In conclusion, modeling the portfolio allocation problem as a structured prediction task over multiple hypotheses allows the separation of the process into two key components: Hypotheses selection and weight optimization. This approach provides parametric methodologies to induce diversity in hypothesis selection through the parameter $\gamma$, and to regulate diversity in the learning process of individual predictors based on the selected hypotheses through the parameter $\varepsilon$. These two sources of diversity are directly connected to out-of-sample portfolio risk diversification and can be controlled and adjusted prior to the optimization and decision-making process.
Figure 3: Asset or hypotheses selection including forecasting (left). 1-moth cumulative return ranking with threshold $T$, randomly selection of $M$ constituents from a sample of $m$ candidates, $\gamma$ is the diversity parameter in hypothesis selection stage (middle). Model architecture for s-RBFN for $m=M$ and $\gamma=1$, and $\varepsilon$ the diversity parameter for the learning stage (right).
4 Numerical Results
4.1 Data and Setup
The time series of daily prices for the S&P 500 index constituents (S&P 500, 2005) over the past six years (excluding weekends) was obtained from Bloomberg (S&P 500, 2024). Returns were calculated as the percentage change in daily prices.
To empirically demonstrate the connection between diversity in the structured prediction setting with multiple hypotheses and out-of-sample portfolio diversification, numerous out-of-sample Sharpe ratios are computed using the predictive setting. The Sharpe ratio is the quotient between return and volatility (risk) for a portfolio of $M$ stocks over a time period $\Delta t$. The Sharpe ratio is computed for different values of portfolio size $M$ and diversity parameter $\varepsilon$, both for the case without diversity in asset selection and for the scenario where the portfolio allocation decision is made once, performing 100 simulations for each trial, as in Modern Portfolio Theory (MPT) (Markowitz, 1952). This allows comparison with traditional methods and reveals the pattern of portfolio diversification behavior in relation to the parameters $M$ and $\varepsilon$.
Next, a more realistic case not considered in MPT is included, where the problem becomes sequential over 10 months, making it more practical and useful for practitioners (in this case, the Sharpe ratio over the 10-month period is reported). For these cases, the asset selection is performed by ranking 1-month cumulative return predictions and choosing the top $M$ assets based on different threshold values $T$. In the one-step decision-making cases, the same assets are usually selected across simulations, whereas in the sequential case, the non-stationarity in the data may alter prediction ranking and asset selection. Finally, a second source of diversity is introduced through asset or hypothesis selection, controlled by the parameter $\gamma$. The experiments are performed for both sources of diversity in one-step as well as in the multi-step decision-making cases.
The structured ensemble models used are the s-RBFN optimized by least squares and two-layer networks as individual predictors with the same tunned hyperparameters for all the cases as described in (Dominguez et al., 2025). The equal-weighted model and the MSE-weighted model are also included, where the weights are either equally distributed across all constituents or inversely proportional to their generalization prediction error. For asset selection, the ranking is constructed by calculating the one-month recurrent prediction of the daily time series and then ranking based on the accumulated monthly return. The prediction is performed using the same model as the individual predictors with identical hyperparameters, but each of the 500 index members has its own network.
4.2 Parametric Portfolio Risk Diversification
Experiments validate the hypothesis that out-of-sample portfolio risk diversification can be controlled prior to the decision making process if the portfolio is modeled as a structured ensemble in a multiple hypothesis prediction setting.
4.2.1 Diversity in the Ensemble Learning Stage
In this section, for each portfolio allocation, asset selection is performed as described previously, taking portfolio constituents as the top $M$ of the ranking based on the 1-month accumulated return that are also higher than the threshold $T$. Two types of experiments are carried out: one-step decision-making (1 month and 100 simulations per trial) and sequential multi-step decision-making (10 months and 1 simulation per trial). The structured prediction model s-RBFN with radial basis and Gaussian functions is used, both without and with regularization. The experiments are repeated for different threshold values $T=-1\%,-0.5\%,0.0\%,0.5\%,1\%$ and diversity parameter $\varepsilon=0,0.1,0.35$.
The result for the one-step decision-making case can be seen in Figure 5, showing a strong similarity with the results from the MPT setup (Markowitz, 1952), which is illustrated in Figure 4. Figure 4 shows the diversification limits according to MPT, where the Sharpe ratios are presented with an inverted axis compared to Figure 5, with the same horizontal axes referring to the number of assets or stocks $M$. In the case of MPT, the Sharpe ratios are shown for different values of the correlation between assets, and it is observed that the negative growth of the correlation has the same effect as the positive growth of the diversity parameter in ensemble learning in the multi-hypothesis prediction setting, as shown in Figure 5. Thus, portfolio diversification, which depends on correlation (the lower, the better) and the number of portfolio constituents, is to some extent equivalent to diversity in the learning of structured ensembles in a setting like the one discussed (regulated by $\varepsilon$) and the number of predictors. The experiments empirically validate the equivalence between out-of-sample portfolio diversification and diversity in the learning of a structured ensemble in a multi-hypothesis prediction setting when used to model a portfolio according to Section 3.2.
Figure 4: The limits of diversification, (Dalio, 2018). Sharpe and Return-to-Risk ratios is the same (inverted axis). The standard deviation is reduced up to a diversification limit.
In Figure 6, the same set of experiments but for the sequential multi-step decision-making case is presented. In this case, the resemblance is different from that shown in MPT and Figure 4, which makes sense since those setups are one-step, whereas the results here are sequential multi-step. Nevertheless, the conclusions remain unchanged, and the same patterns can be observed regarding the relationship between diversity in ensemble learning and out-of-sample portfolio diversification. A consistent increase in Sharpe ratios can be seen as the parameter $\varepsilon$ and the number of predictors, hypotheses, or assets $M$ increase. On the other hand, the regularization parameter has a greater impact in this case than in the one-step scenario shown in Figure 5. It can be observed that the out-of-sample risk-adjusted portfolio performance improves even in the case where the diversity parameter $\varepsilon$ is not applied. This may indicate that performance in multi-step settings depends not only on diversification but also on the complexity of the data. The same experiments can be seen for the case of radial basis functions with the same conclusions in the Appendix in Figures 9 and 10.
(a) $\lambda_{s}=0$; $T=1\%$
(b) $\lambda_{s}=10$; $T=1\%$
(c) $\lambda_{s}=0$; $T=0.5\%$
(d) $\lambda_{s}=10$; $T=0.5\%$
(e) $\lambda_{s}=0$; $T=-0.5\%$
(f) $\lambda_{s}=10$; $T=-0.5\%$
(g) $\lambda_{s}=0$; $T=-1\%$
(h) $\lambda_{s}=10$; $T=-1\%$
Figure 5: One-step (1-month) 100 simulations Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), s-RBFN regularization parameter $\lambda_{s}$, and $-1\%\leq T\leq 1\%$. s-RBFN with Gaussian basis functions.
(a) $\lambda_{s}=0$; $T=1\%$
(b) $\lambda_{s}=10$; $T=1\%$
(c) $\lambda_{s}=0$; $T=0.5\%$
(d) $\lambda_{s}=10$; $T=0.5\%$
(e) $\lambda_{s}=0$; $T=-0.5\%$
(f) $\lambda_{s}=10$; $T=-0.5\%$
(g) $\lambda_{s}=0$; $T=-1\%$
(h) $\lambda_{s}=10$; $T=-1\%$
Figure 6: Multi-step (10-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), s-RBFN regularization parameter $\lambda_{s}$, and $-1\%\leq T\leq 1\%$. s-RBFN with Gaussian basis functions.
4.2.2 Including Diversity in the Asset or Hypothesis Selection Stage
In this section, the same experiments are conducted, but here the $M$ portfolio constituents or hypotheses are randomly selected from the top $m=M\times\gamma$ stocks based on the 1-month cumulative forecast ranking, where $\gamma$, the asset selection diversity parameter, is a multiplier that scales the candidate pool relative to the final selection size $M$. This approach seeks to test the hypothesis that out-of-sample diversification in portfolio allocation is achieved not only during the optimization process—whether through the inclusion of the parameter $\varepsilon$ in this structured prediction model framework or through other plug-in methods without such aid—but also at the stage of hypothesis selection, when choosing assets or stocks prior to optimizing portfolio weights. For that, the experiments are repeated for different values of the multiplier, including the no-diversity case, which coincides with the results in the previous section, with $\gamma=1,2,3,5$ for both one-step-ahead and multi-step-ahead cases, using the s-RBFN model with radial basis functions.
Figures 7 and 8 present the out-of-sample Sharpe ratios for different numbers of stocks $M$, various values of the asset selection diversity parameter $\gamma$, and the learning stage diversity parameter $\varepsilon$, as well as threshold values $T$ used in the asset selection process based on forecast rankings, for the multi-step sequential (10-month strategy) case. Figures 12 and 13 in the Appendix show the same results for the one-step decision case (1-month, 100 simulations).
(a) $\gamma=1;0.5\%$
(b) $\gamma=2;0.5\%$
(c) $\gamma=3;0.5\%$
(d) $\gamma=5;0.5\%$
(e) $\gamma=1;0.0\%$
(f) $\gamma=2;0.0\%$
(g) $\gamma=3;0.0\%$
(h) $\gamma=5;0.0\%$
Figure 7: Multi-step (10-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), multiplicative factor $\gamma$, and $0\%\leq T\leq 0.5\%$. s-RBFN with radial basis functions.
(a) $\gamma=1;-0.5\%$
(b) $\gamma=2;-0.5\%$
(c) $\gamma=3;-0.5\%$
(d) $\gamma=5;-0.5\%$
(e) $\gamma=1;-1\%$
(f) $\gamma=2;-1\%$
(g) $\gamma=3;-1\%$
(h) $\gamma=5;-1\%$
Figure 8: Multi-step (10-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), multiplicative factor $\gamma$, and $-1\%\leq T\leq-0.5\%$. s-RBFN with radial basis functions.
Within the proposed PSEM framework, the Sharpe ratios, along with the average Sharpe ratio for all levels of diversity determined by the structured ensemble diversity parameter $(\varepsilon)$, can also be analyzed to examine the impact of diversity in asset or hypothesis selection on out-of-sample portfolio diversification. These results are presented in Table 1 for the one-step-ahead case (1-month) and Table 2 for the multi-step experiments (10-month strategy), both using the s-RBFN model with radial basis functions.
In the one-step ahead case, a generalized increase in out-of-sample performance, represented by the Sharpe ratio, can be observed when selecting more diverse assets or hypotheses during the stock selection process prior to weight optimization, even if they have worse average predictions. This can be seen in Table 1, where the average Sharpe ratio increases as the values of $m$ increase, driven by increasing value of $\gamma$, allowing for greater diversity of predictors with different out-of-sample performances as part of the stock selection.
This indicates that a higher number of candidates with greater diversity in predictions, and worse average one-month out-of-sample return predictions (with relatively similar RMSE in their predictions), leads to portfolios with better out-of-sample Sharpe ratios and greater diversification, not due to the learning stage, but due to better asset selection. This pattern holds for both thresholds $T=0.5\%$ and $T=0.0\%$, meaning that the variety of 1-month cumulative return predictions from the sample of $m$ candidates includes positive predictions above 0.5% and 0.0%, respectively.
In the case of negative thresholds $T=-0.5\%$ and $T=-1\%$, it can be observed that the pattern is more pronounced, which is surprising as it indicates that adding negative return predictions to the $M$ portfolio constituents benefits portfolio generalization performance and increases out-of-sample diversification. This occurs because it adds diversity to the hypothesis selection, even if it downgrades the performance of the group compared to other selections with more positive predictions but less diversity. This represents the diversity-quality trade-off of return predictions in asset selection discussed in Section 3.4. Just to point out, for $\gamma=2,3,5$ and $M=20,50$, portfolios have constituents selected randomly from samples of $m=40$ to $m=250$, while the index includes 500 stocks, indicating that the spectrum of negative predictions is not small at all.
Table 1 also shows that, as $T$ decreases, increasing diversity improves the average Sharpe ratio up to a limit, with $T=-0.5\%$ yielding better values than $T=-1.0\%$ and the other thresholds for the one-step-ahead case (1-month). This is consistent with the limits of diversification, as shown in Figure 4 (Dalio, 2018).
Table 2 shows similar results, confirming the same hypothesis for sequential multi-step cases (10-month). It is important to note that this scenario is much more restrictive, as it involves making 10 consecutive decisions over 10 months instead of just one. One might expect to prioritize the positiveness of predictions (quality) over the diversity of the prediction set during asset selection in such problems, as this would seem the most logical approach. However, the experiments validate the initial hypothesis and confirm the existence of this trade-off, with a significant impact on generalization performance and out-of-sample risk diversification. In Table 2 it can be observed that with a threshold $T=-0.5\%$ for ranking predictions versus $T=0.5\%$, the average out-of-sample Sharpe ratios are better for experiments with negative $T$ in most cases. Furthermore, in the case of a positive $T$, the results are negative for portfolios with fewer assets, indicating the need for a minimum number of stocks in the portfolio (Table 2, case $T=0.5\%$). This issue can be addressed easily without increasing $M$ by adding diversity in asset or hypothesis selection through $\gamma$. This is very useful for practitioners, if they have constraints in the number of assets in the portfolio, they can increase the level of diversification of their portfolio with the asset selection diversity component.
Finally, it can be shown that both in the one-step ahead case (1-month), which is comparable to the MPT setting and other plug-in methods, and in the multi-step case (10-month), the diversity parameter in the ensemble $(\varepsilon=0,0.1,0.35)$, which relates to weight optimization but not asset selection, shows improvement in all tables for $(\gamma=1)$. Without applying diversity in asset selection and only in optimization, the ratios improve with the parameter and the number of assets, consistent with the results shown in the previous section in more detail.
However, in cases where diversity is included in asset selection prior to optimization with $(\gamma=2,3,5)$, it can be seen that the parameter $\varepsilon$ still has a strong impact, although slightly reduced, on the out-of-sample Sharpe ratio and portfolio diversification. This can only be explained by the fact that diversification results from a combination of diversity in hypothesis selection, given by $\gamma$, and the parametrically introduced diversity in the optimization process through $\varepsilon$ during the learning of individual predictors.
Table 1: One-step-ahead experiments (1 month; 100 simulations): Sharpe ratios for different diversity parameters in the learning stage $\varepsilon=0,0.1,0.35$, number of stocks $M$, and multiplicative factor or diversity parameter in the asset selection stage $\gamma=1,2,3,5$. $T$ represents the threshold in the ranking of return predictions. Avg. Sh. refers to the average Sharpe ratio of the elements in the corresponding column.
Div. $(\varepsilon)$ ($T=0.5\%$) ($T=0.0\%$) ($T=-0.5\%$) ($T=-1\%$)
2 5 10 20 50 2 5 10 20 50 2 5 10 20 50 2 5 10 20 50
Div. $(\gamma=1)$;   $m=M\times\gamma=\{2,5,10,20,50\}$
0 3 5 6 8 10 3 4 6 8 10 4 5 7 8 9 3 5 6 8 10
0.1 4 6 6 9 10 4 7 7 8 10 4 5 8 8 11 1 6 7 8 10
0.35 5 6 7 9 10 5 5 6 8 10 3 6 7 8 10 0 6 8 9 10
Avg. Sh. 4 6 6 8 10 4 5 6 8 10 4 5 7 8 10 1 6 7 8 10
Div. $(\gamma=2)$;   $m=M\times\gamma=\{4,10,20,40,100\}$
0 4 5 6 8 9 4 6 6 7 10 4 5 6 8 10 4 5 5 7 9
0.1 4 5 8 9 11 4 5 7 9 10 4 5 7 8 10 3 5 8 9 11
0.35 3 5 8 9 11 6 8 9 10 10 4 6 7 8 10 3 6 7 9 11
Avg. Sh. 4 5 7 9 11 5 6 7 8 10 4 5 7 8 10 3 5 7 8 10
Div. $(\gamma=3)$;   $m=M\times\gamma=\{6,15,30,60,150\}$
0 4 5 6 7 9 4 6 7 8 9 4 5 6 8 10 5 5 6 8 10
0.1 5 5 7 9 11 6 7 8 11 3 4 6 7 9 1 6 7 8 11
0.35 4 5 7 9 11 6 6 7 9 10 3 5 7 9 11 3 5 7 9 11
Avg. Sh. 4 5 7 9 10 5 6 8 9 7 3 5 7 9 10 3 5 7 9 10
Div. $(\gamma=5)$;   $m=M\times\gamma=\{25,50,100,250\}$
0 6 8 7 11 5 6 8 10 5 7 8 10 5 6 8 8
0.1 6 7 3 6 5 7 9 10 6 8 9 11 5 8 9 12
0.35 5 11 9 11 6 8 10 10 6 8 8 11 6 7 9 11
Avg. Sh 5 9 6 9 5 7 9 10 6 8 8 11 5 7 9 10
Table 2: Multi-step-ahead experiments (10-month sequential strategy): Sharpe ratios for different diversity parameters in the learning stage $\varepsilon=0,0.1,0.35$, number of stocks $M$, and multiplicative factor or diversity parameter in the asset selection stage $\gamma=1,2,3,5$. $T$ represents the threshold in the ranking of return predictions. Avg. Sh. refers to the average Sharpe ratio of the elements in the corresponding column.
Div. $(\varepsilon)$ ($T=0.5\%$) ($T=0.0\%$) ($T=-0.5\%$) ($T=-1\%$)
2 5 10 20 50 2 5 10 20 50 2 5 10 20 50 2 5 10 20 50
Div. $(\gamma=1)$;   $m=M\times\gamma=\{2,5,10,20,50\}$
0 -1 -1 3 8 6 2 5 6 8 10 5 4 9 6 10 5 6 -2 10 11
0.1 0 6 3 15 11 10 14 10 11 11 2 8 1 16 10 1 1 1 14 9
0.35 -3 9 13 10 12 -7 5 8 9 10 8 8 12 6 13 0 -1 0 5 10
Avg. Sh. -1 5 7 11 10 2 8 8 10 10 5 7 7 9 11 2 2 0 10 10
Div. $(\gamma=2)$;   $m=M\times\gamma=\{4,10,20,40,100\}$
0 3 7 12 14 7 10 11 6 4 11 0 5 6 1 11 -4 8 8 7 2
0.1 -1 11 8 9 13 5 3 7 4 11 5 4 5 7 15 -6 10 14 7 11
0.35 5 6 9 10 13 8 8 16 11 10 -2 5 3 9 7 2 13 13 11 12
Avg. Sh. 3 8 10 11 11 8 7 10 7 10 1 4 5 6 11 -3 10 11 8 8
Div. $(\gamma=3)$;   $m=M\times\gamma=\{6,15,30,60,150\}$
0 6 5 10 5 8 -1 2 7 4 10 6 8 0 4 14 4 -2 11 6 10
0.1 11 8 6 12 10 8 11 12 9 11 1 12 5 6 10 1 5 8 8 11
0.35 4 1 0 7 15 8 9 13 9 11 -2 12 8 12 4 -2 11 6 10
Avg. Sh. 7 5 5 8 11 7 7 11 8 11 4 8 8 9 11 -1 3 8 8 10
Div. $(\gamma=5)$;   $m=M\times\gamma=\{25,50,100,250\}$
0 7 9 0 7 6 6 11 11 7 7 8 9 9 4 9 10
0.1 1 19 5 9 7 6 11 13 2 10 11 12 5 11 12 12
0.35 4 13 2 8 8 8 13 11 0 8 8 15 10 3 7 10
Avg. Sh. 4 14 3 8 7 7 11 12 6 6 9 12 8 6 9 11
5 Conclusion and Future Work
This work shows that valuable concepts of modern portfolio theory
(Markowitz, 1952) that are the basis of modern finance can be understood through structured models for multiple hypotheses prediction, providing new insights. By framing the problem this way, the diversity of individual predictors can be linked with portfolio diversification, offering a fresh perspective based on learning in multiple hypothesis prediction settings. The proposed PSEM strategy enables a structured framework for portfolio optimization, using as portfolio target the arithmetic combiner or equal-weighted portfolio. To the authors’ knowledge, the first plug-in method that parametrically controls out-of-sample diversification prior to the decision-making is introduced. This is useful for practitioners as they can access many diversified versions of the structured data set with different levels of diversification before the portfolio optimization.
This work confirms that the out-of-sample diversification and generalization performance of a portfolio depends not only on the optimization process but also on asset selection. To the authors’ knowledge, this is the first time a method that parametrically induces out-of-sample portfolio diversification from the asset selection stage is presented in the literature. Experiments validate the diversity-quality trade-off of return predictions in asset selection discussed in Section 3.4. The experiments show that, for both the one-step and multi-step cases, when the RMSE ranges in prediction are similar, asset selections with greater diversity in the predicted values prevail over less diverse selections, even if the latter have more positive average returns as a group.
Surprisingly, the experiments demonstrate that the out-of-sample diversification can be increased through diversity in asset selection, even with a constant number of constituents (in contrast to diversification limits (Dalio, 2018). This is probably due to the fact that other frameworks do not take into account the diversification component resulting from asset selection diversity. This is particularly useful for practitioners with constraints on the number of assets who aim to diversify their portfolios. The proposed approach offers parametric methods to introduce diversity in the selection of hypotheses via the parameter $\gamma$ and to manage diversity in the training process of individual predictors based on these hypotheses through the parameter $\varepsilon$. Both sources of diversity are closely linked to out-of-sample portfolio risk diversification and can be adjusted prior to the optimization and decision-making stages, which is particularly useful for practitioners.
For future work, it would be interesting to investigate the application of deeper architectures for individual predictors and multimodal datasets. This would enable the construction of portfolios where the hypotheses are not limited to stock time-series but can include any source of information relevant to any particular hypothesis. Additionally, it would be valuable to test the proposed methodology in scenarios where the hypothesis is a group of stocks or a portfolio, and to explore how to choose an appropriate target in such cases. Another potential direction for future research is to examine the relationship between complexity and diversity in the ensemble loss within the PSEM framework. Finally, an interesting application especially in finance would be the creation of a risk premia factor that measures diversity by constructing different indices with varying levels of diversity and monitoring their risk-return profile across all diversity scales.
[…]
Appendix A Diversity in Learning Stage: Radial Basis Functions
(a) $\lambda_{s}=0$; $T=1\%$
(b) $\lambda_{s}=10$; $T=1\%$
(c) $\lambda_{s}=0$; $T=0.5\%$
(d) $\lambda_{s}=10$; $T=0.5\%$
(e) $\lambda_{s}=0$; $T=-0.5\%$
(f) $\lambda_{s}=10$; $T=-0.5\%$
(g) $\lambda_{s}=0$; $T=-1\%$
(h) $\lambda_{s}=10$; $T=-1\%$
Figure 9: One-step (1-month) 100 simulations Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), s-RBFN regularization parameter $\lambda_{s}$, and $-1\%\leq T\leq 1\%$. s-RBFN with radial basis functions.
(a) $\lambda_{s}=0$; $T=1\%$
(b) $\lambda_{s}=10$; $T=1\%$
(c) $\lambda_{s}=0$; $T=0.5\%$
(d) $\lambda_{s}=10$; $T=0.5\%$
(e) $\lambda_{s}=0$; $T=-0.5\%$
(f) $\lambda_{s}=10$; $T=-0.5\%$
(g) $\lambda_{s}=0$; $T=-1\%$
(h) $\lambda_{s}=10$; $T=-1\%$
Figure 10: Multi-step (10-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), s-RBFN regularization parameter $\lambda_{s}$, and $-1\%\leq T\leq 1\%$. s-RBFN with radial basis functions.
Appendix B Ensemble Regularization in the Gaussian Basis Functions Experiments (Multi-step ahead case)
(a) $\varepsilon=0$ ; $\lambda_{s}=0$
(b) $\varepsilon=0.1$ ; $\lambda_{s}=0$
(c) $\varepsilon=0.35$ ; $\lambda_{s}=0$
(d) $\varepsilon=0$ ; $\lambda_{s}=10$
(e) $\varepsilon=0.1$ ; $\lambda_{s}=10$
(f) $\varepsilon=0.35$ ; $\lambda_{s}=10$
Figure 11: Portfolio ensemble 10-month (multi-step) average test RMSE with different diversity parameter $\varepsilon$, ensemble (s-RBFN) regularization parameter $\lambda_{s}$, threshold for asset selection $T$, number of portfolio constituents or predictors (2, 5, 10, 20, 50), and Gaussian basis functions.
To perform a closer analysis of the regularization parameter, it can be seen how for $\lambda_{s}=10$ the generalization error is more stable for all model hyperparameters, and in the case of no regularization parameter, there are some cases in which the results are quite unstable (like in Figures 11(b) and 11(c)) were the test RMSE are displayed. It can be concluded that the regularization parameter reduces the uncertainty of the s-RBFN hyperparameters in the case of the Gaussian basis function. This instability in hyperparameter selection was not observed with the s-RBFN using radial basis functions.
Appendix C Including Diversity in the Asset or Hypothesis Selection Stage: One-step Decision-making
(a) $\gamma=1;0.5\%$
(b) $\gamma=2;0.5\%$
(c) $\gamma=3;0.5\%$
(d) $\gamma=5;0.5\%$
(e) $\gamma=1;0.0\%$
(f) $\gamma=2;0.0\%$
(g) $\gamma=3;0.0\%$
(h) $\gamma=5;0.0\%$
Figure 12: One-step (1-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), multiplicative factor or diversity parameter in the asset selection stage $\gamma$, and $0\%\leq T\leq 0.5\%$. s-RBFN with radial basis functions.
(a) $\gamma=1;T=-0.5\%$
(b) $\gamma=2;T=-0.5\%$
(c) $\gamma=3;T=-0.5\%$
(d) $\gamma=5;T=-0.5\%$
(e) $\gamma=1;T=-1\%$
(f) $\gamma=2;T=-1\%$
(g) $\gamma=3;T=-1\%$
(h) $\gamma=5;T=-1\%$
Figure 13: One-step (1-month) Sharpe ratios (Y-axis) for different number of predictors $M$ (X-axis), diversity parameter $\varepsilon=0,0.1,0.35$ (colours), multiplicative factor or diversity parameter in the asset selection stage $\gamma$, and $-1\%\leq T\leq-0.5\%$. s-RBFN with radial basis functions.